OpenStudy (anonymous):

HELP PLSSS Question: From a collection of 7 books, 4 will be chosen and placed on a shelf arranged from first to fourth. How many different arrangements of 4 books are possible? 1. 24 2. 35 3. 360 4. 840

OpenStudy (kropot72):

The number of permutations is \[7P4=\frac{7!}{(7-4)!}=\frac{7!}{3!}\]